# A Comprehensive Library for Benchmarking Multi-class Visual Anomaly Detection

### v4

@@ Abstract
Visual anomaly detection aims to identify anomalous regions in images through unsupervised learning paradigms, with increasing application demand and value in fields such as industrial inspection and medical lesion detection. Despite significant progress in recent years, there is a lack of comprehensive benchmarks to adequately evaluate the performance of various mainstream methods across different datasets under the practical multi-class setting. The absence of standardized experimental setups can lead to potential biases in training epochs, resolution, and metric results, resulting in erroneous conclusions. This paper addresses this issue by proposing a comprehensive visual anomaly detection benchmark, ADer, which is a modular framework that is highly extensible for new methods. The benchmark includes multiple datasets from industrial and medical domains, implementing fifteen state-of-the-art methods and nine comprehensive metrics. Additionally, we have proposed the GPU-assisted ADEval package to address the slow evaluation problem of metrics like time-consuming mAU-PRO on large-scale data, significantly reducing evaluation time by more than 1000-fold. Through extensive experimental results, we objectively reveal the strengths and weaknesses of different methods and provide insights into the challenges and future directions of multi-class visual anomaly detection. We hope that ADer will become a valuable resource for researchers and practitioners in the field, promoting the development of more robust and generalizable anomaly detection systems. Full codes have been attached in Appendix and will be open-sourced.
@@ 3 Methodology: ADer Benchmark
Method Hyper Params. Efficiency Train Mem. (M) Test Mem. (M) MVTec AD VisA Real-IAD
@@
Aug. DRAEM [37] 8 Adam 1e-4 97.4M 198G UNet 12,602 2,858 3.4h 35s 8.0h 36s 39.2 18m41s
@@
Hybrid UniAD [35] 8 AdamW 1e-4 24.5M 3.4G EN-b4 1,856 968 0.8h 22s 1.0h 18s 4.1h 7m2s
@@ 3.2 VAD Datasets
Dataset Category Number Image Quantity Epoch Setting in ADer
@@ 3.3 Evaluation Metrics
[…]
@@ 3.6 ADEval: Fast and Memory-efficient Routines for mAU-ROC/mAP/mAU-PRO
[…]
@@ 5 Conclusion and Discussion
This paper addresses the urgent need for a comprehensive and fair benchmark in the field of visual anomaly detection. We introduce a modular and scalable ADer library designed to fairly facilitate the evaluation of fifteen advanced VAD methods across multiple mainstream datasets. ensuring a thorough and unbiased assessment of each method’s performance. Our extensive experiments reveal the strengths and weaknesses of different methods, providing valuable insights into their efficiency and training resource consumption. We also develop and open-source a GPU-assisted ADEval package to reduce the evaluation time, enabling extensive assessments. Experimental results highlight the challenges of various VAD methods and offer valuable insights for future research directions.
@@ Appendix B Detailed Quantitative Results on All Dataset Under 100 epochs
Table A1: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3.
@@
Table A2: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. 3.3.
@@
Table A3: Benchmarked results on MVTec LOCO dataset [5] by the suggested metrics in Sec. 3.3.
@@
Table A4: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. 3.3.
@@
Table A5: Benchmarked results on BTAD dataset [28] by the suggested metrics in Sec. 3.3.
@@
Table A6: Benchmarked results on MPDD dataset [20] by the suggested metrics in Sec. 3.3.
@@
Table A7: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. 3.3.
@@
Table A8: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
@@
Table A9: Benchmarked results on Uni-Medical dataset [40] by the suggested metrics in Sec. 3.3.
@@
Table A10: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. 3.3.
@@
Table A11: Benchmarked results on COCO-AD dataset [42] by the suggested metrics in Sec. 3.3.
@@ Appendix C Detailed Quantitative Results on All Dataset Under 300 epochs
Table A12: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3.
@@
Table A13: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. 3.3.
@@
Table A14: Benchmarked results on MVTec LOCO dataset [5] by the suggested metrics in Sec. 3.3.
@@
Table A15: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. 3.3.
@@
Table A16: Benchmarked results on BTAD dataset [28] by the suggested metrics in Sec. 3.3.
@@
Table A17: Benchmarked results on MPDD dataset [20] by the suggested metrics in Sec. 3.3.
@@
Table A18: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. 3.3.
@@
Table A19: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
@@
Table A20: Benchmarked results on Uni-Medical dataset [40] by the suggested metrics in Sec. 3.3.
@@ Appendix D Detailed Quantitative Results on MVTec AD Dataset Under 100 epochs
Table A21: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3.
@@
Table A22: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3.
@@
Table A23: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3.
@@
Table A24: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3.
@@ Appendix E Detailed Quantitative Results on MVTec 3D Dataset Under 100 epochs
Table A25: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. 3.3.
@@
cable_gland Aug. DRAEM [37] 61.1 85.0 89.7 40.1 00.3 00.5 15.7 00.3 41.8
@@
Table A26: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. 3.3.
@@
Table A27: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. 3.3.
@@ Appendix F Detailed Quantitative Results on MVTec LOCO Dataset Under 100 epochs
Table A28: Benchmarked results on MVTec LOCO dataset [5] by the suggested metrics in Sec. 3.3.
@@
juice_bottle Aug. DRAEM [37] 52.1 74.3 83.4 31.3 03.9 10.1 07.3 05.3 37.5
@@
Table A29: Benchmarked results on MVTec LOCO dataset [5] by the suggested metrics in Sec. 3.3.
@@
[…]
splicing_connectors Aug. DRAEM [37] 40.1 54.7 76.4 50.4 06.5 14.1 15.0 07.6 36.7
@@ Appendix G Detailed Quantitative Results on VisA Dataset Under 100 epochs
Table A30: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. 3.3.
@@
Table A31: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. 3.3.
@@
Table A32: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. 3.3.
@@
pipe_fryum Aug. DRAEM [37] 63.5 74.5 83.1 22.9 00.7 02.5 04.9 01.2 36.0
@@
Table A33: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. 3.3.
@@ Appendix H Detailed Quantitative Results on BTAD Dataset Under 100 epochs
Table A34: Benchmarked results on BTAD dataset [28] by the suggested metrics in Sec. 3.3.
@@ Appendix I Detailed Quantitative Results on MPDD Dataset Under 100 epochs
Table A35: Benchmarked results on MPDD dataset [20] by the suggested metrics in Sec. 3.3.
@@
[…]
bracket_black Aug. DRAEM [37] 39.0 51.7 74.6 45.1 00.1 00.4 24.0 00.2 33.6
@@
bracket_brown Aug. DRAEM [37] 42.5 61.1 79.7 46.8 00.8 01.9 27.5 01.0 37.2
@@
bracket_white Aug. DRAEM [37] 34.0 44.3 66.7 42.3 00.0 00.1 09.0 00.0 28.1
@@
Table A36: Benchmarked results on MPDD dataset [20] by the suggested metrics in Sec. 3.3.
@@
[…]
metal_plate Aug. DRAEM [37] 41.9 76.3 84.5 36.3 10.8 21.2 13.3 11.9 40.6
@@
Table A37: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. 3.3.
@@
Table A38: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. 3.3.
@@
Table A39: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. 3.3.
@@
Table A40: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
@@
Table A41: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
@@
Table A42: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
@@
Table A43: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
@@
Table A44: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. 3.3.
@@
Table A45: Benchmarked results on Uni-Medical dataset [40] by the suggested metrics in Sec. 3.3.
@@
Table A46: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. 3.3.
@@
button_battery Aug. DRAEM [37] 34.4 46.9 72.4 51.8 00.5 00.9 17.8 00.4 32.1
@@
Table A47: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. 3.3.
@@
fire_hood Aug. DRAEM [37] 56.3 39.5 53.4 42.7 00.1 00.2 09.9 00.1 28.9
@@
Table A48: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. 3.3.
@@
phone_battery Aug. DRAEM [37] 51.7 46.1 58.0 45.8 00.1 00.2 14.6 00.1 30.9
@@
plastic_plug Aug. DRAEM [37] 64.1 52.9 54.8 41.6 00.0 00.1 09.4 00.0 31.8
@@
Table A49: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. 3.3.
@@
[…]
porcelain_doll Aug. DRAEM [37] 45.1 29.1 49.8 46.6 00.1 00.1 14.6 00.1 26.5
@@
rolled_strip_base Aug. DRAEM [37] 53.3 68.5 79.8 37.1 00.1 00.3 09.4 00.1 35.5
@@
Table A50: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. 3.3.
@@
Table A51: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. 3.3.
@@
Table A52: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. 3.3.
@@
wooden_beads Aug. DRAEM [37] 52.8 44.9 60.2 43.9 00.1 00.3 12.5 00.1 30.7
@@
Table A53: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. 3.3.
@@
Table A54: Benchmarked results on COCO-AD dataset [42] by the suggested metrics in Sec. 3.3.
@@
Table A55: Benchmarked results on COCO-AD dataset [42] by the suggested metrics in Sec. 3.3.
@@
Table A56: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. 3.3.

### v6

@@ Abstract
Visual anomaly detection aims to identify anomalous regions in images through unsupervised learning paradigms, with increasing application demand and value in fields such as industrial inspection and medical lesion detection. Despite significant progress in recent years, there is a lack of comprehensive benchmarks to adequately evaluate the performance of various mainstream methods across different datasets under the practical multi-class setting. The absence of standardized experimental setups can lead to potential biases in training epochs, resolution, and metric results, resulting in erroneous conclusions. This paper addresses this issue by proposing a comprehensive visual anomaly detection benchmark, ADer, which is a modular framework that is highly extensible for new methods. The benchmark includes multiple datasets from industrial and medical domains, implementing fifteen state-of-the-art methods and nine comprehensive metrics. Additionally, we have proposed the GPU-assisted ADEval package to address the slow evaluation problem of metrics like time-consuming mAU-PRO on large-scale data, significantly reducing evaluation time by more than 1000-fold. Through extensive experimental results, we objectively reveal the strengths and weaknesses of different methods and provide insights into the challenges and future directions of multi-class visual anomaly detection. We hope that ADer will become a valuable resource for researchers and practitioners in the field, promoting the development of more robust and generalizable anomaly detection systems.
@@ 3 Methodology: ADer Benchmark
Method Hyper Params. Efficiency
Train Mem.
Test Mem.
MVTec AD VisA Real-IAD
@@
Aug.
DRAEM [37] 8 Adam 1e-4 97.4M 198G UNet 12,602 2,858 3.4h 35s 8.0h 36s 39.2 18m41s
@@
Hybrid
UniAD [35] 8 AdamW 1e-4 24.5M 3.4G EN-b4 1,856 968 0.8h 22s 1.0h 18s 4.1h 7m2s
@@ 3.2 VAD Datasets
Dataset Category Number Image Quantity
Epoch Setting
in ADer
@@ 3.3 Evaluation Metrics
Following the ViTAD [40] setting, we select image-level mean Area Under the Receiver Operating Curve (mAU-ROC) [37], mean Average Precision (mAP) [37], mean $F_{1}$-score ($F_{1}-m$ax) [47], region-level mean Area Under the Per-Region-Overlap (mAU-PRO) [4], pixel-level mAU-ROC, mAP, $F_{1}-m$ax, and the average AD (mAD) [40] of seven metrics to evaluate all experiments. Additionally, we adopt the more practical order-independent pixel-level mean maximal Intersection over Union (mIoU-max) proposed in InvAD [42].
@@ 3.6 ADEval: Fast and Memory-efficient Routines for mAU-ROC/mAP/mAU-PRO
$\begin{aligned} &\widehat{tpr}_{k}=\frac{\left|\{i\mid i\in I\wedge y_{i}=1\wedge\hat{y}_{i}\geq\widehat{thr}_{k}\}\right|}{\left|\{i\mid i\in I\wedge y_{i}=1\}\right|}\\ &=\frac{\sum_{l=k}^{N-1}\left|\{i\mid i\in I\wedge y_{i}=1\wedge\widehat{thr}_{l}\leq\hat{y}_{i}\leq\widehat{thr}_{l+1}\}\right|}{\left|\{i\mid i\in I\wedge y_{i}=1\}\right|},\end{aligned}$
$\begin{aligned} &\widehat{fpr}_{k}=\frac{\left|\{i\mid i\in I\wedge y_{i}=0\wedge\hat{y}_{i}\geq\widehat{thr}_{k}\}\right|}{\left|\{i\mid i\in I\wedge y_{i}=0\}\right|}\\ &=\frac{\sum_{l=k}^{N-1}\left|\{i\mid i\in I\wedge y_{i}=0\wedge\widehat{thr}_{l}\leq\hat{y}_{i}\leq\widehat{thr}_{l+1}\}\right|}{\left|\{i\mid i\in I\wedge y_{i}=0\}\right|},\end{aligned}$
@@ 5 Conclusion and Discussion
This paper addresses the urgent need for a comprehensive and fair benchmark in the field of visual anomaly detection. We introduce a modular and scalable ADer library designed to fairly facilitate the evaluation of fifteen advanced VAD methods across multiple mainstream datasets, ensuring a thorough and unbiased assessment of each method’s performance. Our extensive experiments reveal the strengths and weaknesses of different methods, providing valuable insights into their efficiency and training resource consumption. We also develop and open-source a GPU-assisted ADEval package to reduce the evaluation time, enabling extensive assessments. Experimental results highlight the challenges of various VAD methods and offer valuable insights for future research directions.
@@ Appendix B Detailed Quantitative Results on All Dataset Under 100 epochs
Table A1: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. $3.$
@@
Table A2: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. $3.$
@@
Table A3: Benchmarked results on MVTec LOCO dataset [5] by the suggested metrics in Sec. $3.$
@@
Table A4: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. $3.$
@@
Table A5: Benchmarked results on BTAD dataset [28] by the suggested metrics in Sec. $3.$
@@
Table A6: Benchmarked results on MPDD dataset [20] by the suggested metrics in Sec. $3.$
@@
Table A7: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. $3.$
@@
Table A8: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. $3.$
@@
Table A9: Benchmarked results on Uni-Medical dataset [40] by the suggested metrics in Sec. $3.$
@@
Table A10: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. $3.$
@@
Table A11: Benchmarked results on COCO-AD dataset [42] by the suggested metrics in Sec. $3.$
@@ Appendix C Detailed Quantitative Results on All Dataset Under 300 epochs
Table A12: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. $3.$
@@
Table A13: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. $3.$
@@
Table A14: Benchmarked results on MVTec LOCO dataset [5] by the suggested metrics in Sec. $3.$
@@
Table A15: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. $3.$
@@
Table A16: Benchmarked results on BTAD dataset [28] by the suggested metrics in Sec. $3.$
@@
Table A17: Benchmarked results on MPDD dataset [20] by the suggested metrics in Sec. $3.$
@@
Table A18: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. $3.$
@@
Table A19: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. $3.$
@@
Table A20: Benchmarked results on Uni-Medical dataset [40] by the suggested metrics in Sec. $3.$
@@ Appendix D Detailed Quantitative Results on MVTec AD Dataset Under 100 epochs
Table A21: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. $3.$
@@
Table A22: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. $3.$
@@
Table A23: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. $3.$
@@
Table A24: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. $3.$
@@ Appendix E Detailed Quantitative Results on MVTec 3D Dataset Under 100 epochs
Table A25: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. $3.$
@@
cable_gland
Aug.
DRAEM [37] 61.1 85.0 89.7 40.1 00.3 00.5 15.7 00.3 41.8
@@
Table A26: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. $3.$
@@
Table A27: Benchmarked results on MVTec 3D dataset [6] by the suggested metrics in Sec. $3.$
@@ Appendix F Detailed Quantitative Results on MVTec LOCO Dataset Under 100 epochs
Table A28: Benchmarked results on MVTec LOCO dataset [5] by the suggested metrics in Sec. $3.$
@@
juice_bottle
Aug.
DRAEM [37] 52.1 74.3 83.4 31.3 03.9 10.1 07.3 05.3 37.5
@@
Table A29: Benchmarked results on MVTec LOCO dataset [5] by the suggested metrics in Sec. $3.$
@@
mAU-ROC mAP $F_{1}-m$ax mAU-ROC mAP $F_{1}-m$ax
splicing_connectors
Aug.
DRAEM [37] 40.1 54.7 76.4 50.4 06.5 14.1 15.0 07.6 36.7
@@ Appendix G Detailed Quantitative Results on VisA Dataset Under 100 epochs
Table A30: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. $3.$
@@
Table A31: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. $3.$
@@
Table A32: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. $3.$
@@
pipe_fryum
Aug.
DRAEM [37] 63.5 74.5 83.1 22.9 00.7 02.5 04.9 01.2 36.0
@@
Table A33: Benchmarked results on VisA dataset [47] by the suggested metrics in Sec. $3.$
@@ Appendix H Detailed Quantitative Results on BTAD Dataset Under 100 epochs
Table A34: Benchmarked results on BTAD dataset [28] by the suggested metrics in Sec. $3.$
@@ Appendix I Detailed Quantitative Results on MPDD Dataset Under 100 epochs
Table A35: Benchmarked results on MPDD dataset [20] by the suggested metrics in Sec. $3.$
@@
mAU-ROC mAP $F_{1}-m$ax mAU-ROC mAP $F_{1}-m$ax
bracket_black
Aug.
DRAEM [37] 39.0 51.7 74.6 45.1 00.1 00.4 24.0 00.2 33.6
@@
bracket_brown
Aug.
DRAEM [37] 42.5 61.1 79.7 46.8 00.8 01.9 27.5 01.0 37.2
@@
bracket_white
Aug.
DRAEM [37] 34.0 44.3 66.7 42.3 00.0 00.1 09.0 00.0 28.1
@@
Table A36: Benchmarked results on MPDD dataset [20] by the suggested metrics in Sec. $3.$
@@
mAU-ROC mAP $F_{1}-m$ax mAU-ROC mAP $F_{1}-m$ax
metal_plate
Aug.
DRAEM [37] 41.9 76.3 84.5 36.3 10.8 21.2 13.3 11.9 40.6
@@
Table A37: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. $3.$
@@
Table A38: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. $3.$
@@
Table A39: Benchmarked results on MAD_Real dataset [46] by the suggested metrics in Sec. $3.$
@@
Table A40: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. $3.$
@@
Table A41: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. $3.$
@@
Table A42: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. $3.$
@@
Table A43: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. $3.$
@@
Table A44: Benchmarked results on MAD_Sim dataset [46] by the suggested metrics in Sec. $3.$
@@
Table A45: Benchmarked results on Uni-Medical dataset [40] by the suggested metrics in Sec. $3.$
@@
Table A46: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. $3.$
@@
button_battery
Aug.
DRAEM [37] 34.4 46.9 72.4 51.8 00.5 00.9 17.8 00.4 32.1
@@
Table A47: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. $3.$
@@
fire_hood
Aug.
DRAEM [37] 56.3 39.5 53.4 42.7 00.1 00.2 09.9 00.1 28.9
@@
Table A48: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. $3.$
@@
phone_battery
Aug.
DRAEM [37] 51.7 46.1 58.0 45.8 00.1 00.2 14.6 00.1 30.9
@@
plastic_plug
Aug.
DRAEM [37] 64.1 52.9 54.8 41.6 00.0 00.1 09.4 00.0 31.8
@@
Table A49: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. $3.$
@@
mAU-ROC mAP $F_{1}-m$ax mAU-ROC mAP $F_{1}-m$ax
porcelain_doll
Aug.
DRAEM [37] 45.1 29.1 49.8 46.6 00.1 00.1 14.6 00.1 26.5
@@
rolled_strip_base
Aug.
DRAEM [37] 53.3 68.5 79.8 37.1 00.1 00.3 09.4 00.1 35.5
@@
Table A50: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. $3.$
@@
Table A51: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. $3.$
@@
Table A52: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. $3.$
@@
wooden_beads
Aug.
DRAEM [37] 52.8 44.9 60.2 43.9 00.1 00.3 12.5 00.1 30.7
@@
Table A53: Benchmarked results on Real-IAD dataset [32] by the suggested metrics in Sec. $3.$
@@
Table A54: Benchmarked results on COCO-AD dataset [42] by the suggested metrics in Sec. $3.$
@@
Table A55: Benchmarked results on COCO-AD dataset [42] by the suggested metrics in Sec. $3.$
@@
Table A56: Benchmarked results on MVTec AD dataset [3] by the suggested metrics in Sec. $3.$